# Virtual Multiplex Staining for Histological Images using a Marker-wise Conditioned Diffusion Model

### v1

@@ Abstract
Multiplex imaging is revolutionizing pathology by enabling the simultaneous visualization of multiple biomarkers within tissue samples, providing molecular-level insights that traditional hematoxylin and eosin (H&E) staining cannot provide. However, the complexity and cost of multiplex data acquisition have hindered its widespread adoption. Additionally, most existing large repositories of H&E images lack corresponding multiplex images, limiting opportunities for multimodal analysis. To address these challenges, we leverage recent advances in latent diffusion models (LDMs), which excel at modeling complex data distributions utilizing their powerful priors for fine-tuning to a target domain. In this paper, we introduce a novel framework for virtual multiplex staining that utilizes pretrained LDM parameters to generate multiplex images from H&E images using a conditional diffusion model. Our approach enables marker-by-marker generation by conditioning the diffusion model on each marker, while sharing the same architecture across all markers. To tackle the challenge of varying pixel value distributions across different marker stains and to improve inference speed, we fine-tune the model for single-step sampling, enhancing both color contrast fidelity and inference efficiency through pixel-level loss functions. We validate our framework on two publicly available datasets, notably demonstrating its effectiveness in generating up to 18 different marker types with improved accuracy, a substantial increase over the 2-3 marker types achieved in previous approaches. This validation highlights the potential of our framework, pioneering virtual multiplex staining. Finally, this paper bridges the gap between H&E and multiplex imaging, potentially enabling retrospective studies and large-scale analyses of existing H&E image repositories.
Figure 1: Sample from Orion-CRC (Lin et al. 2023): (a) H&E staining and (b) multiplex immunofluorescence (mIF) of the same region. Each color in mIF corresponds to a different marker.
@@ 1 Introduction
Histopathological analysis is crucial for disease diagnosis and biomedical research, enabling the identification of pathological changes and guiding treatment decisions. Among available techniques, hematoxylin and eosin (H&E) staining has long been the gold standard, often complemented by immunohistochemistry (IHC). Despite their widespread use, both H&E and IHC have limitations in capturing the complex tissue microenvironment, especially in advanced cancer studies (Wharton Jr et al. 2021). To address these limitations, multiplex immunofluorescence (mIF) and multiplex immunohistochemistry (mIHC) imaging have emerged, enabling visualization of multiple biomarkers within a single tissue sample and providing more comprehensive insights into the tissue microenvironment (see Fig. 1(Sood et al. 2016; Tan et al. 2020; Lin et al. 2023). However, widespread adoption of multiplex imaging is hindered by complex imaging protocols and high staining costs.
@@
To overcome these challenges, computational approaches known as virtual staining, which generate various staining images from H&E stained tissue images, have gained much attention (Latonen et al. 2024). Most prior work applied deep generative models to unpaired H&E-to-IHC tasks, focusing on spatial alignment (Li et al. 2023; Chen et al. 2024; Wang et al. 2025); other studies explored multiplex marker generation, such as H&E-to-IF (Burlingame et al. 2020) and H&E-to-mIHC (Pati et al. 2024). However, these approaches generally require a separate model per marker, limiting scalability and increasing computational cost, which hinders efficient handling of large marker sets. Moreover, their separate training manner focuses only on individual markers, missing valuable cross-channel relationships and overlooking the potential benefits of knowledge sharing through joint training. Even recent methods like VIMs (Dubey et al. 2024) and HEMIT (Bian et al. 2024), which generate up to two or three markers, still lack the scalability needed for practical multiplex imaging applications.
@@
Based on these observations, we propose a novel framework for virtual multiplex staining that utilizes a conditional diffusion model (Ho, Jain, and Abbeel 2020; Dhariwal and Nichol 2021; Rombach et al. 2022) to translate H&E images into multiplex marker images. By designing a marker-wise generation approach, we aim to generate multiple markers within a single shared model architecture, enabling knowledge sharing during training and reducing the significant computational overhead. Specifically, to optimize both the efficiency of the generation process and the quality of the generated results, we adopt a two-stage training approach, each addressing a core challenge:
@@
① Multi-target generation. Unlike previous methods training separate models for each marker or lacking scalability for the large number of marker types, we use a marker-wise conditioning strategy to generate multiple markers simultaneously using a single model. By conditioning H&E images with different marker-type embeddings using latent diffusion models (Rombach et al. 2022), we enable the model to effectively distinguish among a large set of markers and generate high-quality multiplex staining from H&E images, overcoming the limitations of conventional text-based conditioning when the number of marker types increases. Therefore, the training strategy reduces computational costs and enables potential knowledge sharing.
@@ 2.1 Multiplex Imaging for Pathology
H&E staining is the gold standard for tissue examination in pathology, with IHC staining complementing morphological assessment by providing additional molecular information. Previous virtual staining studies have mainly focused on translating H&E to IHC images in unpaired settings to address spatial alignment challenges inherent between these stains (Li et al. 2023; Chen et al. 2024; Wang et al. 2025). While these methods have advanced unpaired translation, their objectives and data settings differ from our focus on paired H&E-to-multiplex virtual staining. Moreover, both H&E and IHC stainings are limited in capturing the complex tissue microenvironment, particularly in advanced cancers (Wharton Jr et al. 2021). This has led to the increased interest in multiplex imaging techniques, which simultaneously visualize multiple biomarkers within a single tissue section—enhancing cellular, molecular, and spatial understanding of pathology (Tan et al. 2020; Wharton Jr et al. 2021; Lin et al. 2023). Despite its analytic advantages, the widespread adoption of multiplex imaging is hindered by resource-intensive protocols and limited data availability.
To address these challenges, research has shifted toward deep learning-based virtual staining methods that generate multiplex images from H&E sections (Burlingame et al. 2020; Pati et al. 2024; Bian et al. 2024). Early efforts relied on pix2pix (Isola et al. 2017), a conditional GAN used as a baseline for paired image-to-image translation from H&E to specific immunomarkers and requiring a separate model per marker, thus limiting scalability as the number of biomarkers increases (Burlingame et al. 2020). Recently, HEMIT (Bian et al. 2024) combined residual CNNs with Swin Transformers to jointly generate up to three mIHC markers, partially capturing cross-marker relationships; however, it has yet to demonstrate scalability to larger marker panels. Meanwhile, VIMs (Dubey et al. 2024) used an unpaired, diffusion-based approach to generate two IHC markers with performance comparable to pix2pix, but it requires expert-designed text prompts for marker control, which poses barriers when scaling to more markers or deploying in automated pipelines.
@@ 2.2 Conditional Diffusion Models
Diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020) have emerged as a powerful generative framework, achieving remarkable fidelity in image synthesis via iterative denoising. Of particular interest, Latent Diffusion Models (LDMs) (Rombach et al. 2022) leverage a Variational Autoencoder (VAE) trained on large-scale datasets to operate in a compressed latent space, substantially improving computational efficiency and providing strong priors for rapid adaptation in diverse domains.
Recent studies have demonstrated the versatility of diffusion models for a wide range of image generation. Conditional diffusion models allow fine-grained control over image generation by leveraging diverse conditioning inputs, including text (Saharia et al. 2022), semantic maps (Zhang, Rao, and Agrawala 2023), or images (Tumanyan et al. 2023). Notably, these models have shown strong performance in dense prediction tasks like semantic segmentation, depth estimation, and structured image-to-image translation by effectively integrating conditioning signals into the diffusion process (Zhao et al. 2023; Ke et al. 2024; Fu et al. 2024). Building on these developments, approaches like Marigold (Ke et al. 2024) keep the VAE prior from pre-trained LDMs and fine-tune only the denoising U-Net for specific tasks, such as depth prediction, achieving efficient adaptation with reduced training complexity. Similarly, Geowizard (Fu et al. 2024) introduces a class-wise one-hot conditioning strategy to differentiate multiple geometric and domain classes, which is beneficial when scaling to many output targets conditioned on the same image prompt.
@@ 3.1 Marker-wise Conditional Diffusion Model
For the multiplex image generation, we build our virtual staining model on the pre-trained stable diffusion (SD) backbone (Rombach et al. 2022), leveraging its LDM architecture. Specifically, we formulate the virtual staining as a conditional diffusion process that maps H&E images to generate multiplex marker images.
@@
We replicate a single-channel marker image $\mathbf{I}_{m}$ into three channels $\mathbf{I}^{*}_{m}\in\mathbb{R}^{3\times H\times W}$, where $m$ denotes a marker type from the marker type set $\{1,2,\ldots,M\}$, to match the input shape of the pre-trained SD VAE $\mathcal{E}$. Passing through $\mathcal{E}$, we obtain the latent representation of the H&E image as $\mathbf{x}$ and that of the marker image as $\mathbf{z_{m,0}}$.
@@
Following the DDPM framework (Ho, Jain, and Abbeel 2020), we define a timestep $t\in\{1,...,T\}$ and a variance schedule $\{\beta_{1},...,\beta_{T}\}$ for diffusion process. During training, we add noise to the marker latent representation at timestep $t$ as:
$\mathbf{z}_{m,t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{m,0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$
where $\mathbf{z}_{m,0}$ represent the clean latent embedding and $\mathbf{z}_{m,t}$ denotes the noised version at timestep $t$ for the $m$-th marker type, the notation $\epsilon\sim\mathcal{N}(0,I)$ and $\bar{\alpha}_{t}=\prod_{k=1}^{t}(1-\beta_{k})$. This process gradually adds noise to the latent representation, allowing the model to learn the denoising process.
By freezing the SD VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ parameters, we train only the conditional diffusion U-Net, which serves as the denoising network $\hat{\mathbf{v}}_{m,t}(\cdot)$. Specifically, we concatenate a latent pair $\mathbf{x},\mathbf{z}_{m,t}]$ along the channel dimension at a timestep $t$ as the input. To accommodate this input, we double the U-Net’s input channel size by duplicating weights and halving their values (Ke et al. 2024). The U-Net $\hat{\mathbf{v}}_{m,t}=\hat{\mathbf{v}}_{\theta}([\mathbf{x},\mathbf{z}_{m,t}],t)$ is trained according to the following loss function,
@@
where $\mathbf{v}^{*}_{m,t}=\sqrt{\bar{\alpha}_{t}}\epsilon-\sqrt{(1-\bar{\alpha}_{t})}\mathbf{z}_{m,0}$. This squared error $\mathcal{L}_{m}$ ensures that the model learns to accurately predict the added noise during the forward process, thereby improving its performance during inference. During inference, the SD VAE decoder $\mathcal{D}$ can reconstruct the image from the generated $\hat{\mathbf{z}}_{m,0}$ from noise, i.e., $\hat{\mathbf{I}}_{m}=\mathcal{D}(\mathbf{\hat{z}}_{m,0})$. Notably, the reconstructed image $\hat{\mathbf{I}}_{m}$ is averaged across channels to produce the final prediction, a single-channel marker image $\mathbf{I}_{m}\in\mathbb{R}^{H\times W}$.
@@
Inspired by Fu et al. (2024), we implement marker-wise conditioning by representing each marker type as a one-hot vector. Unlike text conditioning, which can be ambiguous with many marker types, we adopt marker-wise one-hot embedding to provide distinct and scalable signals for multi-marker generation. The one-hot vector undergoes positional encoding and is then element-wise added to the time embedding to condition the model on the specific marker type $m$. This technique enables the diffusion U-Net to efficiently learn marker-specific features, allowing for multiple marker image generation. A detailed analysis is provided in Sec. 4.
@@ 3.2 Fine-tuning for Color Contrast Fidelity
While our model achieves multiplexed image generation in Sec. 3.1, color contrast fidelity remains a primary challenge due to dataset bias toward dark background regions. The visual comparison is detailed in Sec. 4. Moreover, efficient processing can also be important for large-scale pathology data, as diffusion models are intrinsically slow due to their iterative process. While approaches such as DDIM (Song, Meng, and Ermon 2020) help accelerate inference, other techniques like test-time ensembling (Ke et al. 2024) introduce computational overhead for improved performance.
To address both color fidelity and inference efficiency, we introduce a second, fine-tuning stage as shown in Fig. 2(c). In this stage, inspired by recent latent diffusion advances (Garcia et al. 2024), we enable single-step inference by fixing $t=T$ and replacing the random noise with zero noise ($\epsilon=0$) , optimizing the model directly in pixel space. This modification allows for rapid, deterministic image generation and, crucially, supports the direct application of task-specific pixel-level supervision for improved color contrast. Consistent with previous training phases, we update only the parameters of the diffusion U-Net, while keeping the VAE encoder and decoder fixed.

### v2

@@ Abstract
Multiplex imaging is revolutionizing pathology by enabling the simultaneous visualization of multiple biomarkers within tissue samples, providing molecular-level insights that traditional hematoxylin and eosin (H&E) staining cannot provide. However, the complexity and cost of multiplex data acquisition have hindered its widespread adoption. Additionally, most existing large repositories of H&E images lack corresponding multiplex images, limiting opportunities for multimodal analysis. To address these challenges, we leverage recent advances in latent diffusion models (LDMs), which excel at modeling complex data distributions by utilizing their powerful priors for fine-tuning to a target domain. In this paper, we introduce a novel framework for virtual multiplex staining that utilizes pretrained LDM parameters to generate multiplex images from H&E images using a conditional diffusion model. Our approach enables marker-by-marker generation by conditioning the diffusion model on each marker, while sharing the same architecture across all markers. To tackle the challenge of varying pixel value distributions across different marker stains and to improve inference speed, we fine-tune the model for single-step sampling, enhancing both color contrast fidelity and inference efficiency through pixel-level loss functions. We validate our framework on two publicly available datasets, notably demonstrating its effectiveness in generating up to 18 different marker types with improved accuracy, a substantial increase over the 2-3 marker types achieved in previous approaches. This validation highlights the potential of our framework, pioneering virtual multiplex staining. Finally, this paper bridges the gap between H&E and multiplex imaging, potentially enabling retrospective studies and large-scale analyses of existing H&E image repositories.
Figure 1: Sample from Orion-CRC (lin2023high): (a) H&E staining and (b) multiplex immunofluorescence (mIF) of the same region. Each color in mIF corresponds to a different marker.
@@ 1 Introduction
Histopathological analysis is crucial for disease diagnosis and biomedical research, enabling the identification of pathological changes and guiding treatment decisions. Among available techniques, hematoxylin and eosin (H&E) staining has long been the gold standard, often complemented by immunohistochemistry (IHC). Despite their widespread use, both H&E and IHC have limitations in capturing the complex tissue microenvironment, especially in advanced cancer studies (wharton2021tissue). To address these limitations, multiplex immunofluorescence (mIF) and multiplex immunohistochemistry (mIHC) imaging have emerged, enabling visualization of multiple biomarkers within a single tissue sample and providing more comprehensive insights into the tissue microenvironment (see Fig. 1(sood2016multiplexed; tan2020overview; lin2023high). However, widespread adoption of multiplex imaging is hindered by complex imaging protocols and high staining costs.
@@
To overcome these challenges, computational approaches known as virtual staining, which generate various staining images from H&E stained tissue images, have gained much attention (latonen2024virtual). Most prior work applied deep generative models to unpaired H&E-to-IHC tasks, focusing on spatial alignment (li2023adaptive; chen2024pathological; wang2025oda); other studies explored multiplex marker generation, such as H&E-to-IF (burlingame2020shift) and H&E-to-mIHC (pati2024accelerating). However, these approaches generally require a separate model per marker, limiting scalability and increasing computational cost, which hinders efficient handling of large marker sets. Moreover, their separate training manner focuses only on individual markers, missing valuable cross-channel relationships and overlooking the potential benefits of knowledge sharing through joint training. Even recent methods like VIMs (dubey2024vims) and HEMIT (bian2024hemit), which generate up to two or three markers, still lack the scalability needed for practical multiplex imaging applications.
@@
Based on these observations, we propose a novel framework for virtual multiplex staining that utilizes a conditional diffusion model (ho2020denoising; dhariwal2021diffusion; rombach2022high) to translate H&E images into multiplex marker images. By designing a marker-wise generation approach, we aim to generate multiple markers within a single shared model architecture, enabling knowledge sharing during training and reducing the significant computational overhead. Specifically, to optimize both the efficiency of the generation process and the quality of the generated results, we adopt a two-stage training approach, each addressing a core challenge:
@@
① Multi-target generation. Unlike previous methods training separate models for each marker or lacking scalability for the large number of marker types, we use a marker-wise conditioning strategy to generate multiple markers simultaneously using a single model. By conditioning H&E images with different marker-type embeddings using latent diffusion models (rombach2022high), we enable the model to effectively distinguish among a large set of markers and generate high-quality multiplex staining from H&E images, overcoming the limitations of conventional text-based conditioning when the number of marker types increases. Therefore, the training strategy reduces computational costs and enables potential knowledge sharing.
@@ 2.1 Multiplex Imaging for Pathology
H&E staining is the gold standard for tissue examination in pathology, with IHC staining complementing morphological assessment by providing additional molecular information. Previous virtual staining studies have mainly focused on translating H&E to IHC images in unpaired settings to address spatial alignment challenges inherent between these stains (li2023adaptive; chen2024pathological; wang2025oda). While these methods have advanced unpaired translation, their objectives and data settings differ from our focus on paired H&E-to-multiplex virtual staining. Moreover, both H&E and IHC stainings are limited in capturing the complex tissue microenvironment, particularly in advanced cancers (wharton2021tissue). This has led to the increased interest in multiplex imaging techniques, which simultaneously visualize multiple biomarkers within a single tissue section—enhancing cellular, molecular, and spatial understanding of pathology (tan2020overview; wharton2021tissue; lin2023high). Despite its analytic advantages, the widespread adoption of multiplex imaging is hindered by resource-intensive protocols and limited data availability.
To address these challenges, research has shifted toward deep learning-based virtual staining methods that generate multiplex images from H&E sections (burlingame2020shift; pati2024accelerating; bian2024hemit). Early efforts relied on pix2pix (isola2017image), a conditional GAN used as a baseline for paired image-to-image translation from H&E to specific immunomarkers, which required a separate model per marker, thus limiting scalability as the number of biomarkers increases (burlingame2020shift). Recently, HEMIT (bian2024hemit) combined residual CNNs with Swin Transformers to jointly generate up to three mIHC markers, partially capturing cross-marker relationships; however, it has yet to demonstrate scalability to larger marker panels. Meanwhile, VIMs (dubey2024vims) used an unpaired, diffusion-based approach to generate two IHC markers with performance comparable to pix2pix, but it requires expert-designed text prompts for marker control, which poses barriers when scaling to more markers or deploying in automated pipelines.
@@ 2.2 Conditional Diffusion Models
Diffusion models (sohl2015deep; ho2020denoising) have emerged as a powerful generative framework, achieving remarkable fidelity in image synthesis via iterative denoising. Of particular interest, Latent Diffusion Models (LDMs) (rombach2022high) leverage a Variational Autoencoder (VAE) trained on large-scale datasets to operate in a compressed latent space, substantially improving computational efficiency and providing strong priors for rapid adaptation in diverse domains.
Recent studies have demonstrated the versatility of diffusion models for a wide range of image generation. Conditional diffusion models allow fine-grained control over image generation by leveraging diverse conditioning inputs, including text (saharia2022photorealistic), semantic maps (zhang2023adding), or images (tumanyan2023plug). Notably, these models have shown strong performance in dense prediction tasks like semantic segmentation, depth estimation, and structured image-to-image translation by effectively integrating conditioning signals into the diffusion process (zhao2023unleashing; ke2024repurposing; fu2025geowizard). Building on these developments, approaches like Marigold (ke2024repurposing) keep the VAE prior from pre-trained LDMs and fine-tune only the denoising U-Net for specific tasks, such as depth prediction, achieving efficient adaptation with reduced training complexity. Similarly, Geowizard (fu2025geowizard) introduces a class-wise one-hot conditioning strategy to differentiate multiple geometric and domain classes, which is beneficial when scaling to many output targets conditioned on the same image prompt.
@@ 3.1 Marker-wise Conditional Diffusion Model
For the multiplex image generation, we build our virtual staining model on the pre-trained stable diffusion (SD) backbone (rombach2022high), leveraging its LDM architecture. Specifically, we formulate the virtual staining as a conditional diffusion process that maps H&E images to generate multiplex marker images.
@@
We replicate a single-channel marker image $\mathbf{I}_{m}$ into three channels to match the input shape of the pre-trained SD VAE $\mathcal{E}$, where $m\in\{1,2,...,M\}$ denotes a marker type. During training, we encode the ground-truth marker image through the VAE encoder to obtain the clean latent representation $\mathbf{z}_{m,0}=\mathcal{E}(\mathbf{I}_{m})$. Correspondingly, the H&E image is encoded as $\mathbf{x}$.
@@
Following the DDPM framework (ho2020denoising), we define a timestep $t\in\{1,...,T\}$ and a variance schedule $\{\beta_{1},...,\beta_{T}\}$ for diffusion process. Specifically, during the training stage, we corrupt the ground-truth marker latent $\mathbf{z}_{m,0}$ by adding noise at timestep $t$:
$\mathbf{z}_{m,t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{m,0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (1)
where $\epsilon\sim\mathcal{N}(0,I)$ is Gaussian noise and $\bar{\alpha}_{t}=\prod_{k=1}^{t}(1-\beta_{k})$. Note that $\mathbf{z}_{m,0}$ is only available during training from ground-truth marker images. This forward process gradually adds noise to the latent representation, allowing the U-Net to learn the denoising operation.
By freezing the SD VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ parameters, we train only the conditional diffusion U-Net, which serves as the denoising network $\hat{\mathbf{v}}_{m,t}(\cdot)$. Specifically, we concatenate a latent pair $\mathbf{x},\mathbf{z}_{m,t}]$ along the channel dimension at a timestep $t$ as the input. To accommodate this input, we double the U-Net’s input channel size by duplicating weights and halving their values (ke2024repurposing). We train only the U-Net denoising network $\hat{\mathbf{v}}_{m,t}=\hat{\mathbf{v}}_{\theta}([\mathbf{x},\mathbf{z}_{m,t}],t)$ using the following loss function:
@@
where $\mathbf{v}^{*}_{m,t}=\sqrt{\bar{\alpha}_{t}}\epsilon-\sqrt{(1-\bar{\alpha}_{t})}\mathbf{z}_{m,0}$ represents the training target for noise prediction. This formulation ensures the model learns to accurately predict the noise added during the forward diffusion process.
Inference procedure.
During inference, the generation process starts from random noise $\mathbf{z}_{m,T}\sim\mathcal{N}(0,I)$ and iteratively denoises using the learned U-Net. Starting from $t=T$ and stepping down to $t=1$, we sample:
$\mathbf{z}_{m,t-1}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathbf{z}_{m,t}-\sqrt{1-\bar{\alpha}_{t}}\hat{\mathbf{v}}_{m,t}([\mathbf{x},\mathbf{z}_{m,t}],t,c_{m}).$ (3)
Note that we do not use ground-truth $\mathbf{z}_{m,0}$ during inference. After reaching $t=0$, the final denoised latent $\hat{\mathbf{z}}_{m,0}$ is reconstructed into pixel space via the frozen VAE decoder: $\mathbf{\hat{I}}_{m}=\mathcal{D}(\hat{\mathbf{z}}_{m,0})$. The reconstructed image is averaged across channels to produce a single-channel marker image prediction.
@@
Inspired by fu2025geowizard, we implement marker-wise conditioning by representing each marker type as a one-hot vector. Unlike text conditioning, which can be ambiguous with many marker types, we adopt marker-wise one-hot embedding to provide distinct and scalable signals for multi-marker generation. The one-hot vector undergoes positional encoding and is then element-wise added to the time embedding to condition the model on the specific marker type $m$. This technique enables the diffusion U-Net to efficiently learn marker-specific features, allowing for multiple marker image generation. A detailed analysis is provided in Sec. 4.
@@ 3.2 Fine-tuning for Color Contrast Fidelity
While our model achieves multiplexed image generation in Sec. 3.1, color contrast fidelity remains a primary challenge due to dataset bias toward dark background regions. The visual comparison is detailed in Sec. 4. Moreover, efficient processing can also be important for large-scale pathology data, as diffusion models are intrinsically slow due to their iterative process. While approaches such as DDIM (song2020denoising) help accelerate inference, other techniques like test-time ensembling (ke2024repurposing) introduce computational overhead for improved performance.
To address both color fidelity and inference efficiency, we introduce a second, fine-tuning stage as shown in Fig. 2(c). In this stage, inspired by recent latent diffusion advances (garcia2024fine), we enable single-step inference by fixing $t=T$ and replacing the random noise with zero noise ($\epsilon=0$) , optimizing the model directly in pixel space. This modification allows for rapid, deterministic image generation and, crucially, supports the direct application of task-specific pixel-level supervision for improved color contrast. Consistent with previous training phases, we update only the parameters of the diffusion U-Net, while keeping the VAE encoder and decoder fixed.